OpenStudy (anonymous):

divide 5x^2+10 by x^4-4. what is it in simpliest terms

OpenStudy (anonymous):

5x^2 + 10 = 5(x^2 + 2). x^4 - 4 = (x^2 + 2)(x^2 - 2) So, the division gives: 5/(x^2-2)